archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    1 bor Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details The 10G card crashed was restarted The ppp session ant the l2tp peer will restart on the spare lns Date Tuesday 17

    Original URL path: http://status.ovh.es/?do=details&id=3077 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details We had an electrical interruption on the NRA The equipment have rebooted We verify with France Telecom to find out what there was Date Friday 20 July

    Original URL path: http://status.ovh.es/?do=details&id=3085 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    us to secure the data exchange between France Telecom and OVH has expired France Telecom works urgently to send us a new certificate Date Tuesday 24 July 2012 12 38PM Reason for closing Done Additional comments about closing The problem is solved The eligibility test is operational again Comments 1 Comment by OVH Tuesday 24 July 2012 00 37AM Certificates have just been placed The SSL connection is performing as

    Original URL path: http://status.ovh.es/?do=details&id=3114 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    anonymous invalid tag 27806 JN 009c1faa 73ee232e1 To sip XXXXXXX XXXXXX user phone tag 13a67a0 6dbe2c01 13c4 50029 27bcd1 2dbeeea3 27bcd1 Call ID 27806 AF 009c1fa9 60c29bfc5 sip ovh net CSeq 10171414 INVITE Via SIP 2 0 UDP XXXXXXXXX 5962 branch z9hG4bK NLVP 25b7c40c 44721ac1 Supported replaces 100rel Content Length 0 Date Wednesday 25 July 2012 13 03PM Reason for closing Done Comments 3 Comment by OVH Wednesday 25 July

    Original URL path: http://status.ovh.es/?do=details&id=3128 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Incident Category Backend Core Status CLOSED Percent Complete Details Worktasks on the optical fiber connecting the sites of BRIGNAIS to VENISSIEUX are in progress Securing the optical loop is playing its role perfectly flowing traffic of subscribers to the backbone routers Subscribers are not impacted Date Friday 03 August 2012 01 34AM Reason for closing Done Additional comments about closing Worktask is done now All sites of loop

    Original URL path: http://status.ovh.es/?do=details&id=3163 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    firewall cards tonight http status ovh net do details id 3105 some modems cannnot login Date Thursday 09 August 2012 01 17AM Reason for closing Done Comments 1 Comment by OVH Wednesday 08 August 2012 14 27PM We took a

    Original URL path: http://status.ovh.es/?do=details&id=3185 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    unbundled customers We cut the links between the unbundling infrastructure and of this router until identifying the origin of the problem DHCP on Lyon works normally by now Date Friday 10 August 2012 16 41PM Reason for closing Done Comments 3 Comment by OVH Friday 10 August 2012 02 01AM We are going to update the switches managing the xDSL infrastructure of Lyon We will begin with lyo 3 n5

    Original URL path: http://status.ovh.es/?do=details&id=3206 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the optical fiber linking the Prado sites to Sainte Marguerite are processing securing the optical buckle plays perfectly its role by flowing the traffic of subscribers to the backbone routers Subscribers are not impacted Date Tuesday 21 August 2012 02

    Original URL path: http://status.ovh.es/?do=details&id=3252 (2015-10-15)
    Open archived version from archive